OpenStudy (anonymous):

Determine the discriminant of x2 = 4. 17 16 -15 -10

OpenStudy (anonymous):

discriminant D=b^2-4ac

rishavraj (rishavraj):

ax^2 + bx + c = 0 D = b^2 - 4ac in question b = 0

OpenStudy (anonymous):

in ur question b is 0. a is 1 and c is 4

rishavraj (rishavraj):

@conain007 hey c = -4

OpenStudy (anonymous):

yes c=-4

rishavraj (rishavraj):

@roseromeroo u got it??

OpenStudy (anonymous):

@rishavraj @conain007 i'm gonna try it out now thank you both :)

rishavraj (rishavraj):

u r most welcome:))

OpenStudy (anonymous):

yw

OpenStudy (andrewhaze):

are you new @roseromeroo ?

OpenStudy (anonymous):

@andrewhaze yes haha

OpenStudy (andrewhaze):

welcom to OS then :)

OpenStudy (andrewhaze):

and if you ever need help just tag me :P

OpenStudy (anonymous):

@andrewhaze okay! thank you :)